OpenStudy (anonymous):

A total of 478 tickets were sold for the school play. They were either adult tickets or student tickets. There were 72 fewer student tickets sold than adult tickets. How many adult tickets were sold? Sorry - apparently need this last one for this topic.

OpenStudy (johnweldon1993):

so s + a = 478 we know that s = a -72 plug that into the original equation a - 72 + a = 478 add 72 to both sides a + a = 550 2a = 550 a = 275 so there were 275 adult tickets so 275 + s = 478 s = 203 is 203 ...72 fewer than 275? yes it is indeed so those are your answers s = 203 and a = 275 and yes s = student a = adult

OpenStudy (anonymous):

Thank you!

OpenStudy (johnweldon1993):

no problem....but do you understand the logic behind it? setting up 2 equations...solving for 1....plugging it into the other..solving for the other...and checking?

OpenStudy (anonymous):

Yes. :)